# Experimental End-to-End Optimization of Directly Modulated Laser-based IM/DD Transmission

### v1

@@ IV-B Benchmark RX-side equalization schemes
The performance of the E2E scheme is compared to two (FFE, VNLE) RX-side equalization benchmark schemes, giving context on the performance gains obtained by using E2E learning. As in the case of the E2E scheme, the transmission link is based on the experimental setup in Fig. 2b. The DSP structure of the benchmark schemes is represented in Fig. 5. The TX structure of both FFE and VNLE benchmark schemes is identical: both feature an equispaced 4PAM mapper, 2-SpS upsampling, and a 17-tap RRC cosine filter with roll-off factor $\alpha=0.1$. The obtained discrete sequence is then linearly upsampled to 65 GSa/s for digital-to-analog conversion. Given that $P_{\mathrm{RF}}$ is not numerically optimized in the case of the benchmark setups, $P_{\mathrm{RF}}$ is swept in the range [4,2][-4,2] dBm. $I_{\mathrm{bias}}$ is set to the same value as the one learned by the E2E scheme to allow a fair comparison. On the RX side, the signal is first downsampled to 2 SpS. The two benchmark setups differ only in the RX equalization method. While the linear equalization scheme uses a 20-tap FIR FFE, the nonlinear one is based on a 2nd order VNLE, using 20 taps for the 1st order kernel and 10 taps for the nonlinear kernel. The 20 linear taps correspond to the total length of the pulse shaping and EQ filters of the E2E setup. By employing both linear and nonlinear equalization as benchmarks, we quantify the impact of nonlinearities on the link’s performance across $R_{s}$ and $L$. After equalization, the signal is downsampled to 1 SpS and detected using 1-tap maximum likelihood detection (MLD). Given that the MLD provides hard decision thresholding, the SER is calculated as the fraction of detection errors to the total number of symbols $N$.
@@ V Experimental transmission results
The performance comparison of the E2E and RX EQ benchmark schemes includes 2 different baudrates (20 and 30 GBd) and 3 different transmission distances (0, 1, 2 km) for a total of 6 cases. The hyperparameters in the optimization of the E2E scheme (learning rate, number of FIR filter taps at TX and RX) are simultaneously fine-tuned using multi-dimensional grid search. The reported results correspond to the scheme achieving minimum SER when tested on the experimental setup. This process is conducted for every $R_{s}$ and $L$ independently. When calculated from the experimental setup, the SER is calculated based on a sequence of $10^{6}$ symbols. The RX EQ benchmarks are trained offline using experimentally obtained sequences. The reported SER results are obtained using previously unseen sequences for a total length of $10^{6}$ symbols. The experimentally obtained SER results for $R_{s}=\{20,\;30\}$ GBd and $L=\{0,\;1,\;2\}$ km with respect to $P_{\mathrm{RF}}$ are plotted in Fig. 7.
@@
Looking at the $L=1$ km case, where the CD and chirp interaction starts to have an impact on the performance, the SER performance difference between schemes stay relatively the same. The E2E scheme achieves the lower SER again with respect to the VNLE ($2.44\times 10^{-4}$ vs $1.36\times 10^{-3}$ at 20 GBd and $8.75\times 10^{-3}$ vs $1.13\times 10^{-2}$ at 30 GBd). In absolute terms, the 20 GBd error rates remain similar to B2B transmission, while the 30 GBd ones are about an order of magnitude higher for all non-DB schemes. This advantage may be due to the fact that low-baudrate signals are more resilient to CD [38]. The DB-based schemes show virtually no performance difference with the B2B case at both $R_{s}$, thanks to their 50% smaller signal bandwidth.
@@
Looking at the $L=0$ km case, it becomes apparent that the eye opening of the E2E signal is larger than its counterpart. Firstly, the E2E scheme’s signal shows a larger distance between adjacent levels than the RRC-shaped signal. Additionally, the extinction ratio between the highest and lowest power symbols is also higher. Given that both signals are subject to the same power constraints at the TX side, the advantage of the E2E scheme over the RRC-shaped signal must come from the constellation and pulse shaping. This can also be seen in the eye width of the pulses in the E2E signal compared to the RRC-shaped one. While the RRC-shaped signal decays rapidly toward the center of the eye, the E2E signal decays noticeably slower. Another interesting difference is the eye skew in both signals. The symbols in the RRC-shaped signal are clearly not horizontally aligned, despite the practical absence of CD. This is likely due to gain compression in the laser cavity, that leads to some symbols having a longer rise time than others. This effect is mostly absent in the E2E signal, which could also explain the lower error rate in the latter. A similar trend can be seen in the 1 and 2 km transmission cases. The chirp-CD interaction can induce power fading in both signals, reducing their SNR at the RX side. However, they are not impacted in the same way. The RRC-shaped signals suffers severe eye closure at 1 km and its eye becomes indistinguishable at 2 km. On the other hand, the impact of the higher transmission distance on the E2E signal is significantly lower. While the signal is attenuated as the transmission distance increases, its 4 levels remain distinguishable even for $L=2$ km, explaining the SER results in Fig. 7. The advantage of the E2E scheme points again toward an efficient level tuning and pulse shape design of the E2E signal with respect to the RRC-shaped signal, allowing it to deliver better performance in the presence of CD.

### v2

@@ IV-B Benchmark RX-side equalization schemes
The performance of the E2E scheme is compared to two (FFE, VNLE) RX-side equalization benchmark schemes, giving context on the performance gains obtained by using E2E learning. As in the case of the E2E scheme, the transmission link is based on the experimental setup in Fig. 3b. The DSP structure of the benchmark schemes is represented in Fig. 6. The TX structure of both FFE and VNLE benchmark schemes is identical: both feature an equispaced 4PAM mapper, 2-SpS upsampling, and a 17-tap RRC cosine filter with roll-off factor $\alpha=0.1$. The obtained discrete sequence is then linearly upsampled to 65 GSa/s for digital-to-analog conversion. Given that $P_{\mathrm{RF}}$ is not numerically optimized in the case of the benchmark setups, $P_{\mathrm{RF}}$ is swept in the range [4,2][-4,2] dBm. $I_{\mathrm{bias}}$ is set to the same value as the one learned by the E2E scheme to allow a fair comparison. On the RX side, the signal is first downsampled to 2 SpS. The two benchmark setups differ only in the RX equalization method. While the linear equalization scheme uses a 20-tap FIR FFE, the nonlinear one is based on a 2nd order VNLE, using 20 taps for the 1st order kernel and 10 taps for the nonlinear kernel. The 20 linear taps correspond to the total length of the pulse shaping and EQ filters of the E2E setup. The FFE-based structure of the E2E setup gives it equivalent $\mathcal{O}(N)$ complexity to the FFE benchmark setup. By employing both linear and nonlinear equalization as benchmarks, we quantify the impact of nonlinearities on the link’s performance across $R_{s}$ and $L$. After equalization, the signal is downsampled to 1 SpS and detected using 1-tap maximum likelihood detection (MLD). Given that the MLD provides hard decision thresholding, the SER is calculated as the fraction of detection errors to the total number of symbols $N$.
@@ V Experimental transmission results
The performance comparison of the E2E and RX EQ benchmark schemes includes 2 different baudrates (20 and 30 GBd) and 3 different transmission distances (0, 1, 2 km) for a total of 6 cases. The baudrates are chosen to approximately match the bandwidth of the laser at low $I_{\mathrm{bias}}$. The hyperparameters in the optimization of the E2E scheme (learning rate, number of FIR filter taps at TX and RX) are simultaneously fine-tuned using multi-dimensional grid search. The reported results correspond to the scheme achieving minimum SER when tested on the experimental setup. This process is conducted for every $R_{s}$ and $L$ independently. When calculated from the experimental setup, the SER is calculated based on a sequence of $10^{6}$ symbols. The RX EQ benchmarks are trained offline using experimentally obtained sequences. The reported SER results are obtained using previously unseen sequences for a total length of $10^{6}$ symbols. The experimentally obtained SER results for $R_{s}=\{20,\;30\}$ GBd and $L=\{0,\;1,\;2\}$ km with respect to $P_{\mathrm{RF}}$ are plotted in Fig. 7.
@@
Looking at the $L=1$ km case, where the CD and chirp interaction starts to have an impact on the performance, the SER performance difference between schemes stay relatively the same. The E2E scheme achieves the lower SER again with respect to the VNLE ($2.44\times 10^{-4}$ vs $1.36\times 10^{-3}$ at 20 GBd and $8.75\times 10^{-3}$ vs $1.13\times 10^{-2}$ at 30 GBd). In absolute terms, the 20 GBd error rates remain similar to B2B transmission, while the 30 GBd ones are about an order of magnitude higher for all non-DB schemes. This advantage may be due to the fact that low-baudrate signals are more resilient to CD [Zhang:18]. The DB-based schemes show virtually no performance difference with the B2B case at both $R_{s}$, thanks to their 50% smaller signal bandwidth.
@@
Looking at the $L=0$ km case, it becomes apparent that the eye opening of the E2E signal is larger than its counterpart. Firstly, the E2E scheme’s signal shows a larger distance between adjacent levels than the RRC-shaped signal. Additionally, the extinction ratio between the highest and lowest power symbols is also higher. Given that both signals are subject to the same power constraints at the TX side, the advantage of the E2E scheme over the RRC-shaped signal must come from the constellation and pulse shaping. This can also be seen in the eye width of the pulses in the E2E signal compared to the RRC-shaped one. While the RRC-shaped signal decays rapidly toward the center of the eye, the E2E signal decays noticeably slower. Another interesting difference is the eye skew in both signals. The symbols in the RRC-shaped signal are clearly not horizontally aligned, despite the practical absence of CD. This is likely due to gain compression in the laser cavity, that leads to some symbols having a longer rise time than others. This effect is mostly absent in the E2E signal, which could also explain the lower error rate in the latter. A similar trend can be seen in the 1 and 2 km transmission cases. The chirp-CD interaction can induce power fading in both signals, reducing their SNR at the RX side. However, they are not impacted in the same way. The RRC-shaped signals suffers severe eye closure at 1 km and its eye becomes indistinguishable at 2 km. On the other hand, the impact of the higher transmission distance on the E2E signal is significantly lower. While the signal is attenuated as the transmission distance increases, its 4 levels remain distinguishable even for $L=2$ km, explaining the SER results in Fig. 7. The difference in amplitude between the three eye diagrams in Fig. 8 (top) also shows the substantial power fading introduced by the interaction between CD and chirp as $L$ is increased, which is less pronounced for the E2E scheme. The advantage of the E2E scheme points again toward an efficient level tuning and pulse shape design of the E2E signal with respect to the RRC-shaped signal, allowing it to deliver better performance in the presence of CD.